Pls Sir solve if a,b,c are in GP and logc base a,Loga base b,logb base c are in ap find common  difference of ap

Grade:12th pass

1 Answers

Arun
5 years ago
a, b, c are in GP => b2 = ac -------(1)
Also log (ac) + log(ab) = 2log(bc)
  • Log(a2bc) = log (b2c2)
  • a2bc = b2c2
  • a= bc ------(2)
from 1 and 2 , a=b      
common difference of AP = log(ac/bc) = log(a/b) = log(1)
= 0


Regards
Arun